OpenStudy (anonymous):

(10 5/8+3 5/8)+ 2 3/8=n+12 Solve for n

OpenStudy (anonymous):

It's for my little sister :)

OpenStudy (anonymous):

37/8

OpenStudy (anonymous):

thats not an option :/

OpenStudy (anonymous):

I need the actual answer.

OpenStudy (anonymous):

what are the options

OpenStudy (anonymous):

A. 6 5/6 B. 18 5/8 C. 4 5/8 D. 2 5/8

OpenStudy (anonymous):

@freakenpurplepizza

OpenStudy (anonymous):

(10 5/8 + 3 5/8) + 2 3/8 = n + 12 (45/8 + 29/8) + 19/8 = n + 12 (multiply by 8) 45 + 29 + 19 = 8n + 96 93 = 8n + 96 93 - 96 = 8n - 3 = 8n -3/8 = n

OpenStudy (anonymous):

Thanks :)